three men can complete a certain work in 7 days. after one day's work if three more men join, the remaining work would be completed in how many days
Answer by ankor@dixie-net.com(22740)

three men can complete a certain work in 7 days.
after one day's work if three more men join, the remaining work would be completed in how many days
:
Find the number of man-days to complete the job
3 * 7 = 21 man-days
:
"after one day's work if three more men join, the remaining work would be completed in how many days"
:
let d = number of additional days
3(d+1) + 3d = 21
3d + 3 + 3d = 21
6d = 21 - 3
d = 18/6
d = 3 additional days to complete the job